OpenStudy (anonymous):

suppose that the greatest common divisor of the positive integers X and Y is 21 and the least common multiple of X and Y is 252. The sum of all possible values of X is: a) 20 b) 28 c)420 d)588

OpenStudy (anonymous):

the greatest common divisor is the same as the greatest common factor btw.

myininaya (myininaya):

so we need to look at \[\gcd(x,y)=21 ; lcm(x,y)=\frac{x \cdot y}{\gcd(x,y)}=\frac{x \cdot y}{21} =252 \] \[x \cdot y =252(21)\]

myininaya (myininaya):

so we can play with these x's and y's right here

myininaya (myininaya):

so here are two examples of what x can be 252,21 remember we need x>21

myininaya (myininaya):

126,42

myininaya (myininaya):

63,84

myininaya (myininaya):

i think that is it x can be 252,21,126,42,63,84

myininaya (myininaya):

so the sum of these is 588

myininaya (myininaya):

all of those pairs i wrote have gcd 21

myininaya (myininaya):

guess we should also check to see if they have lcm 252

myininaya (myininaya):

we know the first pair does

myininaya (myininaya):

the second pair does not 126=126*1 126=42*3 lcm(42,126)=126 so we should throw that pair out

myininaya (myininaya):

so we have 252,21,63,84 sum is 420 but what is the lcm of 63 and 84...

myininaya (myininaya):

ok lcm(63,84) is 252 my bad so possible x's are 252,21,63,84 the sum of those is 420